OpenStudy (anonymous):

find the equation that describes the set of points (x,y,z) that are 5 units from the point (2,-1,6)

zepdrix (zepdrix):

Hmmmm can you imagine what shape this function will take? :) If we have a point.. and our function is 5 units from this point in every direction. :x We could think of 5 as our *cough* radius *cough*!

OpenStudy (anonymous):

a circle!! :O

zepdrix (zepdrix):

Good good :) But in 3-dimenionssssssssss it's a ....?

OpenStudy (anonymous):

sphere

OpenStudy (anonymous):

so do you use the same standard equation but with another for the third point?

zepdrix (zepdrix):

Yes we start with the standard equation,\[\Large x^2+y^2+z^2=r^2\] It looks like in this case 5 will be our r. And we need to `shift` x,y, and z based on the `center point` that they gave us.

OpenStudy (anonymous):

(x+2)^2 + (x+1)^2 + (x-6)^2 =5^2

zepdrix (zepdrix):

Hmm not sure why all of those are x's :) lol

OpenStudy (anonymous):

oh yeah theyd be x,y,z

zepdrix (zepdrix):

And your first shift looks a little off, maybe that was a typo. If our point is at x=2, our first bracket should be,\[\large (x-2)^2\]

OpenStudy (anonymous):

oh yeah that was a typo :) THANKS!!!!!!!!!!